Recent Log Entries

Report of SL06 HVPS on with cooling fan off

Log Number: 3814897Logbook(s): Entry Makers: eformanBody: 

Received report from facilities of SL06 HVPS on with cooling fan off (the vein switch should interlock against this condition). EES-RF on-call contacted, is checking it out.

Operator SIGN IN

Log Number: 3814896Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Mock
DATE:             07/15/2020
SHIFT:               Swing

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

1L12 frequency measurement after pumpdown

Log Number: 3814893Logbook(s): Entry Makers: gciovatiBody: 

A network analyzer is presently connected to cavity 7 in 1L12 to measure the resonant frequency after pumpdown to 2.07 K. REF and PROBE cables are also disconnected from the control module for the same reason.
The resonant frequencies of all 8 cavities will be measured, then all cables will be reconnected to the control modules. Larry King is helping with this task.

[PD Update] Evening Plans

Log Number: 3814892Logbook(s): Tags: Body: 

Resweep N. Linac, Do injector studies once NLinac is in proper state, recover RF skipping 1L12 until it is filled, no LEMMING, push cavities to DRVHs, turn off and submit OPS-PR for problem children to be addressed tomorrow, GSET=0 for 1L25-8. Wear masks appropriately at all times.

iocnl07rffcc7 restoration failed

Log Number: 3814891Logbook(s): Tags: Body: restorelib 2.1 Target restore date was 2020-07-15 10:05:36 The output from the myRestore execution was as follows: iocnl07rffcc7 Rf12Fcc Pause: 60 second(s) Execute: /cs/opshome/burt/REBOOT/Rf12Fcc/DLOAD_ICAL.csh iocnl07rffcc7 (success) c100_rffcc@2020-07-15 10:05 The following have no download value R177WFSCOPrun R177WFTRIPrun 32(34) restored Execute: burtwb -r 3 -f /cs/opshome/burt/REBOOT/Rf12Fcc/snap-ecal/iocnl07rffcc7-ecal.snap (success) Execute: burtwb -r 3 -f /cs/opshome/burt/REBOOT/Rf12Fcc/snap-exp/iocnl07rffcc7-exp.snap (s

NL Dropped to Indeterminate State

Log Number: 3814889Logbook(s): Entry Makers: bcadeBody: 

NL dropped to indeterminate state from controlled access. I was able to safely process two individuals into the North Linac. However, when they were leaving, the door locks would not disengage. One person was actually stuck in the personnel access trap. There was nothing unusual seen under the section labeled 'North Access Room.' Switch A, B to the right of the Exch Key were both labeled red and the release indicator (square) was red. This was normal because one of the keys were not returned.

CREW CHIEF SIGN IN

Log Number: 3814888Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 07-15-20
Time : 15:39:42
Day : Wednesday
Shift : Swing

Crew Chief : Forman
Program Deputy: McGuckin
Team : X
Arm : RADCON
Designated SSO: Mock
LERF Operator:
Operators : Mock

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

1L07-7 FCC

Log Number: 3814885Logbook(s): Body: 

Replaced the FCC chassis due to com issues. MD

Follow-up Re: 1L15-1 Klystron Problem

Log Number: 3814884Logbook(s): Body: 

Mod anode bd. replaced by D. Gelhaar and I to fix bogus KMAV. Then a RFCM and I/O buffer card to fix no gradient problem. Cavity back to RF at ODVH=11.5 Mv/m.

IOC MAC address change

Log Number: 3814878Logbook(s): Tags: Body: Event/action ID: 3 [mod.type.prop.ioc.macaddress]
Deployment: ced
Workspace: ops
Changes:
iocnl07rffcc7.macaddress

Force Open NL VBV1L11B/VBV1L12B

Log Number: 3814877Logbook(s): Entry Makers: marbleBody: 

~ 12:30 pm VBV1L11B and VBV1L12B were forced open and Limit switch for VBV1L13A was reset. All Good

MQA4A12 Mismatch

Log Number: 3814873Logbook(s): Entry Makers: bcadeBody: 

mismatch alarm came in for this magnet. A demand setpoint of the magnet and a channel reset was attempted to clear the mismatch. Neither attempt worked.

Pages